OpenStudy (nuccioreggie):

There are (713)3 ⋅ 70 strawberries in a field. What is the total number of strawberries in the field? 1 716 717 739

OpenStudy (nuccioreggie):

@pooja195

OpenStudy (nuccioreggie):

@Michele_Laino

OpenStudy (michele_laino):

is your equation like this: \[\Large {44^x} = 432\]?

OpenStudy (nuccioreggie):

Im not on that question anymore please help with this one

OpenStudy (michele_laino):

I know, I'm referring to the previous question

OpenStudy (nuccioreggie):

I know but no it wasnt like that please help with this question

OpenStudy (michele_laino):

we can write this: \[\Large {\left( {{7^{13}}} \right)^3} \times {7^0} = {7^{13 \times 3}} \times {7^0} = {7^{13 \times 3 + 0}} = ...\]

OpenStudy (nuccioreggie):

So its D

OpenStudy (nuccioreggie):

??

OpenStudy (nuccioreggie):

@Michele_Laino

OpenStudy (michele_laino):

that's right!